Problem 9: Given X is a random variable with normal distribution, N (?, ?2). Given ?-6 and ?2-0.25, determine b, c, and d such that: I. P (Xs b) - 0.05 Il. P (X2 c) 0.40

i need this step by step please
Answer #1

Given : X\to N(\mu=6,\sigma^2=0.25 )

Therefore , \sigma=\sqrt{0.25}=0.5

I )

P(X\leq b)=0.05

P(X\leq -b)=0.05

P(\frac{X-\mu}{\sigma}\leq -\frac{b-6}{0.5})=0.05

P(Z\leq -\frac{b-6}{0.5})=0.05

From statistical table ,

\therefore -\frac{b-6}{0.5}=1.64

\therefore -b=1.64*0.5-6=-5.18

\therefore b=5.18

II )

P(X\geq c)=0.40

P(\frac{X-\mu}{\sigma}\geq \frac{c-6}{0.5})=0.40

P(Z\geq \frac{c-6}{0.5})=0.40

From statistical table ,

\therefore \frac{c-6}{0.5}=0.25

\therefore c=0.25*0.5+6=6.1250

III )

P(|X-6|\leq d)=0.95

P(-d<X-6<d)=0.95

P(-\frac{d}{0.5}<\frac{X-6}{0.5}<\frac{d}{0.5})=0.95

P(-\frac{d}{0.5}<Z<\frac{d}{0.5})=0.95

P(Z>-\frac{d}{0.5})-P(Z>\frac{d}{0.5})=0.95

1-P(Z<-\frac{d}{0.5})-P(Z>\frac{d}{0.5})=0.95

1-P(Z>\frac{d}{0.5})-P(Z>\frac{d}{0.5})=0.95

1-2*P(Z>\frac{d}{0.5})=0.95

2*P(Z>\frac{d}{0.5})=1-0.95=0.05

P(Z>\frac{d}{0.5})=0.05/2=0.025

From statistical table ,

\therefore \frac{d}{0.5}=1.96

\therefore d=1.96*0.5=0.98